OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation (x+4)^2/9 - (x-5)^2/16 - 1

OpenStudy (anonymous):

\[\frac{ (x+4)^{2} }{ 9 }-\frac{(x-5)^{2}}{16-1}\] Right?

OpenStudy (anonymous):

minus one on the whole thing

OpenStudy (anonymous):

So... \[\frac{ (x+4)^{2} }{ 9 }-\frac{(x-5)^{2}}{16}-1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait, the x and y need to be flipped

OpenStudy (anonymous):

Wait you just need a y

OpenStudy (anonymous):

which side

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[\frac{ (y+4)^{2} }{ 9 }-\frac{(x-5)^{2}}{16}-1\]

OpenStudy (anonymous):

Oh, I swear I read left

OpenStudy (anonymous):

\[\frac{ (x+4)^{2} }{ 9 }-\frac{(y-5)^{2}}{16}-1\]

OpenStudy (anonymous):

\[\frac{ (x+4)^2 }{ 9 } - \frac{ (y-5)^2 }{ 16 } = 1\]

OpenStudy (anonymous):

= 1?

OpenStudy (anonymous):

Yes it's a hyperbola equation

OpenStudy (anonymous):

Thought it was -1

OpenStudy (anonymous):

Mk lemme think for a moment

OpenStudy (anonymous):

x2/a2 − y2/b2 = 1 That's the equation for a hyperbola I think... right?

OpenStudy (anonymous):

Think so

OpenStudy (anonymous):

K I'm reading a refresher so I can remember how to do this

OpenStudy (anonymous):

Okay(:

OpenStudy (anonymous):

Okay, so here's what I understand so far: The hyperbola is centered on (h,k) So our center point is (4,-5) Still reading...

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

We need to use: \[y = \pm \frac{ b }{ a }(x-h) + k\]

OpenStudy (anonymous):

Wait I think I got the answer!!

OpenStudy (anonymous):

Good, because I just realized I haven't learned this yet and I'm struggling.

OpenStudy (anonymous):

I do have the answer though, so I can tell you if it's true or not.

OpenStudy (anonymous):

It would be the Vertices: (-1,5),(-7.5); Foci: (-9,5),(1,5)

OpenStudy (anonymous):

Yup. Can you explain to me how you got it or do you not have time

OpenStudy (anonymous):

I'll just ask a new question so someone can explain it. :D

OpenStudy (anonymous):

Well it's a hyperbola that opens left/right so the center is at (h,k) and to find the vertices you use: (h+-a,k), for foci you use (h+-c,k). I was confused on reading it but you helped me!! Oh and the standard equation for that type of hyperbola is (x-h)^2/a^2 - (y-k)^2/b^2 = 1